Question 3 A monopolist can produce at a constant average and marginal cost of ATC- MC demand demand curve given by Q-53-P. $Hello, could you solve Question3 - Part 3 (the third question) please, Thank you very much!
Answer #1

3) From part 2) we have a demand function P = 53 - (Q1 + Q2)

Profit for firm 1 = π1 = Revenue - costs = PQ1 - 5Q1

π1 = (53 - Q1 - Q2)Q1 - 5Q1

= 53Q1 - Q1^2 - Q1Q2 - 5Q1

π1 = 48Q1 - Q1^2 - Q1Q2

Now profit is maximum when marginal profit is 0

π1 '(Q1) = 0

48 - 2Q1 - Q2 = 0

Q1 = 24 - 0.5Q2

This is the reaction function for firm 1

Similarly we have

Profit for firm 2 = π2 = Revenue - costs = PQ2 - 5Q2

π2 = (53 - Q1 - Q2)Q2 - 5Q2

= 53Q2 - Q1Q2 - Q2^2 - 5Q2

π2 = 48Q2 - Q2^2 - Q1Q2

Now profit is maximum when marginal profit is 0

π2 '(Q2) = 0

48 - 2Q2 - Q1 = 0

Q2 = 24 - 0.5Q1

This is the reaction function for firm 2